Exercise 211 (Part Level Submission) Flint Company reports the following for the month of June. Units Unit Cost Total Cost $9Please show ALL work clearly. Neat handwriting please :)
Answer #1

Total units available for sale = 225+525+750 = 1500 units

Total cost of goods available for sale = 900+2625+4500 = $8025

Average cost per unit = 8025 / 1500 = $5.35 per unit

(a)

FIFO

Ending inventory = 280*6 = $1680

Cost of goods sold = 8025 - 1680 = $6345

LIFO

Ending inventory = 225*4 + (280-225)*5 = $1175

Cost of goods sold = 8025 - 1175 = $6850

Average Cost

Ending inventory = 280*5.35 = $1498

Cost of goods sold = 8025 - 1498 = $6527